OpenStudy (anonymous):

A Bus is Moving at 22m/s [E] for 12s. Then the bus driver slows down at 1.2m/s^2 [W] until the bus stops. Determine the total displacement of the bus.

OpenStudy (anonymous):

You gotta break down the problem and then any problem will seem easy! So lets draw out what this may look like and list our known's and unknown's. We will begin by separating the total displacement into [E] and [W]. To find the displacement for [E] it is fairly simple \[distance=speed \times time \] distance=22m/s x 12seconds= 264m now the second part is a bit harder; we need to find the time first which is the initial speed (22m/s) devided by the rate it took to stop (1.2m/s^2) and we get 18.333 seconds. we put this time into the equation s=1/2at^2 to get 201.66 m. we add up the total displacement and we get 201.666+264=465.66 We also just derived the equation; \[s=ut+1/2at ^{2}\]